3. Trading in foreign exchange

What are spot rates and forward rates?

Suppose you open the newspaper today and observe the following indirect exchange rate quotations for the British pound:

Spot Exchange Rates

Forward Exchange Rates

30 Days

60 Days

90 Days

British pound (pound/dollar) 0.5298 0.5311 0.5335 0.5378

The British pound is selling at a in the forward market.

Suppose you make a £450,000 sale to a British customer who has 60 days to pay you in cash. The customer will pay you in British pounds, but your company is based in the United States, so you are most concerned with the dollar value of the payment. If the customer pays you £450,000 today, how much is that worth in dollars?

$806,908

$637,033

$891,846

$849,377

Assume that the forward market is correct and the 60-day forward exchange rate quoted in the newspaper today (above) is the spot exchange rate 60 days from now. If the customer waits the full 60 days and pays you £450,000, how much have you lost (in dollar terms) due to exchange rate fluctuations?

$5,891

$6,480

$4,713

$6,186
Answer #1

Calculating the worth of amount received from customer today in dollars

Since the customer pays in ful today, therefore

Sale = Amount received from customer in Pounds today = £450,000

Current spot rate of £ per $ = £0.5298 / $

Worth of amount received from customer today in dollars = Amount received from customer in Pounds today / Current spot rate of £ per $ = 450000 / 0.5298 = $ 849377.12 = $849377 (rounded to nearest whole dollar)

Answer $849377

Calculating the loss in dollar terms due exchange rate fluctuations due to delayed payment

Since the customer pays in full in 60 days, therefore

Sale = Amount received from customer in Pounds in 60 days = £450000

Forward rate of £ per $ = Spot rate in 60 days of £ per $ = 0.5335

Worth of amount received from customer after in days in dollars = Amount received from customer in Pounds in 60 days / Spot rate in 60 days of £ per $ = 450000 / 0.5335 = $843486.41 = $843486 (rounded to nearest whole dollar)

Loss = Worth of amount received from customer today in dollars - Worth of amount received from customer in 60 days in dollars = 849377 - 843486 = $5891

Hence Loss due to exchange rate fluctuations = $5891

Answer: $5891